OpenStudy (anonymous):

Two triangles can be formed with the given information. Use the Law of Sines to solve the triangles. B = 29°, b = 26, c = 28 A = 92.5°, C = 58.5°, a = 53.6; A = 87.5°, C = 121.5°, a = 53.6 A = 92.5°, C = 58.5°, a = 12.6; A = 87.5°, C = 121.5°, a = 12.6 A = 119.5°, C = 31.5°, a = 46.7; A = 2.5°, C = 148.5°, a = 2.3 A = 119.5°, C = 31.5°, a = 14.5; A = 2.5°, C = 148.5°, a = 14.5

OpenStudy (anonymous):

Set up the equation for law of sines, sin29/26 = sinC/28 to find the angle of C.

OpenStudy (anonymous):

After you find the angle of C, you can subtract the angles of C and B by 180 to find the angle of A.

OpenStudy (anonymous):

okay give me a sec

OpenStudy (anonymous):

Once you find the angle of A then use law of sines again to find the side of angle A by plugging it in.

OpenStudy (anonymous):

im so lost lol

OpenStudy (anonymous):

Alright give me a few seconds.

OpenStudy (anonymous):

First triangle.

OpenStudy (anonymous):



OpenStudy (anonymous):

We want to now find the angle of c since we are given it's side. So let's set up our equation. sin29/26 = sinc/28 When we solve we get 31.47 for the angle of c. Now we subtract that number by 180 and the angle of b to find the angle of a. So angle a = 119.5.

OpenStudy (anonymous):



OpenStudy (anonymous):

Now we use law of sines again to find the side of a. We can use either the side and angle of B or C to solve for A. But since we know both the angle and side of B let's just use that instead. sin119.52/sidea = sin29/26 So our answer for the side of a is 46.

OpenStudy (anonymous):

Therefore; angle C = 31.47 angle A = 119.5 side a = 46

OpenStudy (anonymous):

theres no choice for a=46

OpenStudy (anonymous):

A = 119.5°, C = 31.5°, a = 46.7; A = 2.5°, C = 148.5°, a = 2.3

OpenStudy (anonymous):

46.67656571 is what I got on my calculator.

OpenStudy (anonymous):

oh okay thank you so much !!!!!

OpenStudy (anonymous):

Anytime.